OpenStudy (australopithecus):

20 packages are processed through customs each hour, let X denote the number of packages that need to be checked every hour. a) if the percentage of packages that need to be checked is 1% then what is the probability that hour 10 is the first sample at which X exceeds 1 Ok, so I calculated, P(X = 0) and P( X = 1) P(X = 0) = 0~Binom(0.1,20) = 0.8179 P(X = 1) = 1~Binom(0.1,20) = 0.1636 So, P(X>1) = 0.0185 Let Y be the number of packages that need to be checked every 10 hours. I'm looking for P(Y>1) = 1 - (P(Y=1) + P(Y=0)) P(Y=0) = (0.8179)^(10) P(Y=1) = (0.8179)^9(0.1636)

OpenStudy (australopithecus):

So I get, P(Y<1) = 1-(0.134 + 0.02679) = 0.83921

OpenStudy (australopithecus):

Is this correct? or am I way off?

OpenStudy (anonymous):

Well, you're not really factoring in what hour 10 means. So I don't quite get that.

OpenStudy (anonymous):

Does hour 10 means that \(10\times 20\) packages have gone through?

OpenStudy (australopithecus):

yes

OpenStudy (australopithecus):

but I assume we have to look at each hour as an individual trial

OpenStudy (australopithecus):

The answer I formulated doesnt really make sense to me though

OpenStudy (australopithecus):

But it sort of does :\

OpenStudy (australopithecus):

My reasoning is that the order of successes and failures in a geometric series doesnt matter as long as there are a certain number of successes and a certain number of successes

OpenStudy (australopithecus):

but I could be wrong

OpenStudy (anonymous):

Is says hour 10 is the first sample. That means that the previous hours got \(X=0\), doesn't it?

OpenStudy (anonymous):

oh wait, now I see...

OpenStudy (australopithecus):

It says when it exceeds 1

OpenStudy (australopithecus):

when X exceeds 1 meats X>1

OpenStudy (anonymous):

so each one before got 0 or 1.

OpenStudy (australopithecus):

means*

OpenStudy (australopithecus):

this question is confusing me so much :\

OpenStudy (anonymous):

Then you want to find: \[ [1-\Pr(X> 1)]^9\Pr(X> 1) \]

OpenStudy (australopithecus):

the thing that confuses me is that a X = 1 in any of the trials, but X has to equal 2 or more in trial 10

OpenStudy (anonymous):

If we treat each hour as being independent.

OpenStudy (australopithecus):

true, but cant that result in X ending up larger than 1 in one of the 9 trials because they are all independent

OpenStudy (australopithecus):

or am I way off maybe I'm not thinking correctly about this it is so hot in my apartment

OpenStudy (anonymous):

Consider the simple case where the first hour doesn't exceed and the second one does.

OpenStudy (anonymous):

Okay, well it is summer.

OpenStudy (australopithecus):

True enough I guess complaining about weather is not something someone should do unless it is destroying their life

OpenStudy (anonymous):

Your concern is that the formula I gave would allow say.... pass exceed pass pass pass pass pass pass pass for example?

OpenStudy (australopithecus):

yes

OpenStudy (australopithecus):

Maybe I dont have a proper understanding of geometric distribution which is probably the case

OpenStudy (australopithecus):

I guess it doesnt matter if that is the case

OpenStudy (australopithecus):

because X>2

OpenStudy (australopithecus):

or equal

OpenStudy (anonymous):

Well, what I would say is that: pass exceed pass pass pass pass pass pass pass Has the same probability as: pass pass pass pass pass pass pass pass exceed If we think of these events as independent. And the only way to allow every combination is if we multiplied by \(9 \choose 1\).

OpenStudy (anonymous):

Or \(10 \choose 1\) I guess.

OpenStudy (anonymous):

So the only real question is whether or not the events are independent.

OpenStudy (anonymous):

If they aren't independent, then where does that \(1\%\) come from?

OpenStudy (australopithecus):

They are independent

OpenStudy (australopithecus):

I guess what you propose makes sense, I may just need to sleep on it I feel like I'm not understanding why it is the right answer.

OpenStudy (anonymous):

Do you remember the binomial distribution formula?

OpenStudy (australopithecus):

yes, \[\left(\begin{matrix}n \\ x\end{matrix}\right)(1-p)^x(p)^{n-x}\]

OpenStudy (anonymous):

Do you understand how it was derived?

OpenStudy (australopithecus):

I think I kind of understand it but perhaps I should reread the chapter on binomials, so I guess no.

OpenStudy (anonymous):

It's very simple. Here, let's come up with a simple example. Consider you want to roll a die 5 times, and you want to get 1 exactly two times.

OpenStudy (anonymous):

So in this case \(p=1/6\), \(n=5\), \(x=2\).

OpenStudy (anonymous):

If we did this manually, we would start with: \[ \frac 16 \times \frac 16\times \frac 56\times \frac 56\times \frac 56 \]This means we got 1 on the first two attempts.

OpenStudy (anonymous):

Now suppose we get 1 on the last two attempts: \[ \frac 56\times \frac 56\times \frac 56\times \frac 16\times \frac 16 \]Due to commutative property of multiplication, these are equivalent.

OpenStudy (anonymous):

So we get the:\[ p^{x}(1-p)^{n-x} \]part here. And it just becomes a question of where do we distribute the the successes?

OpenStudy (anonymous):

And that number comes from \(n\choose x\).

OpenStudy (anonymous):

Does that make sense?

OpenStudy (australopithecus):

yes that does thanks

OpenStudy (anonymous):

You are saying since the probability that it happens at the start is equal to the probability that happens at the end, then somehow we are calculating the probability of either happening.

OpenStudy (anonymous):

When in fact we need to put a \(10\choose 1\) in front if we want to calculate that.

OpenStudy (anonymous):

In which case, the geometric series becomes a binomial distribution with \(k=1\).

OpenStudy (australopithecus):

Ok that makes sense :), thank you for straightening this out for me

OpenStudy (anonymous):

Now one thing to consider... I think is...

OpenStudy (anonymous):

What is the whole Y part about?

OpenStudy (anonymous):

Seem to me that we let \(n=20\times 10\)?

OpenStudy (australopithecus):

Yeah but each trial is independent

OpenStudy (anonymous):

But in this case, is it okay if you get 2 in the very first trial.

OpenStudy (anonymous):

Y is still a binomial bariable at the end of the day, isn't it?

OpenStudy (australopithecus):

It is in the geometric chapter questions in my textbook which comes before poisson so I dont think it can be treated as a poisson distribution

OpenStudy (australopithecus):

The question is basically just saying that in the first 9 hours X=1 for one of the trials and X=0 for the rest of the trials but on the 10th hour X = 2 or more

OpenStudy (anonymous):

Poisson is not very helpful here because we don't know what \(\lambda\) is, do we?

OpenStudy (australopithecus):

so that is where I'm conflicted with your answer wio, because more than one trial in your solution can be X=1, thus X=2 can happen anytime during the 10 trials, but I assume order isnt important so maybe that isnt a factor?

OpenStudy (anonymous):

My answer gives you the probability that it happens on the first trial and not in the last 9. If give you the probability it happens on the last trial but not the first 9. It doesn't give you the probability that either of these events happen.

OpenStudy (anonymous):

For it to do that, you'd have to multiply by 2.

OpenStudy (anonymous):

Wait, do you mean my \(n=10\times 20\) thing?

OpenStudy (australopithecus):

ugh ok ok I understand ughhhhh I'm being dumb sorry

OpenStudy (australopithecus):

maybe you are right perhaps you could just take, 1 - (0~biom(0.1,200) + 1~biom(0.1,200) ) = P(Y>1)

OpenStudy (anonymous):

Ah, biom distribution.

OpenStudy (australopithecus):

I should check to see if it gives me the same answer I got 0.0156 for the one way already went over

OpenStudy (australopithecus):

got a different answer :\

OpenStudy (anonymous):

Why should they give the same answer?

OpenStudy (anonymous):

If you checked 2 packages in the first hour, you would not be counted for the first case, but you would for the second case.

OpenStudy (australopithecus):

ok ok I think I get it thank you so much for your help :)

OpenStudy (kropot72):

Using the binomial distribution I calculated the probability that 2 or more packages need checking in an hourly check of 20 to be 0.0184. Then using the geometric distribution the probability that 2 or more packages need checking for the first time at hour10 is: \[p(10)=0.0184\times0.9816^{10-1}=0.0156\]

OpenStudy (australopithecus):

that is also what I got Kropot72

OpenStudy (kropot72):

Yes, I had noticed that and though it was worth posting. Sorry about posting so many misguided attempts at this (I have deleted them).